OpenStudy (anonymous):

2ln10-ln5 as a single logarithm?

jimthompson5910 (jim_thompson5910):

Any ideas on where to start?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

Are you familiar with any logarithm properties?

OpenStudy (anonymous):

kind of

jimthompson5910 (jim_thompson5910):

Are you familiar with the idea that y*ln(x) can be rewritten as ln(x^y) ?

OpenStudy (anonymous):

okay, yeah sure

jimthompson5910 (jim_thompson5910):

what about the idea that ln(x) - ln(y) = ln(x/y) ?

OpenStudy (anonymous):

hmmm, yes

jimthompson5910 (jim_thompson5910):

that's great

jimthompson5910 (jim_thompson5910):

so 2ln(10) = ln(10^2) using the first property I wrote out

jimthompson5910 (jim_thompson5910):

ln(10^2) = ln(100) So 2ln(10) - ln(5) becomes ln(100) - ln(5) with me so far?

OpenStudy (anonymous):

yes sir

jimthompson5910 (jim_thompson5910):

now use the second property to get ln(100) - ln(5) = ln(100/5) = ln(20) So ln(100) - ln(5) = ln(20) which means 2ln(10) - ln(5) = ln(20)

OpenStudy (anonymous):

oooooohhhhh, okay, makes sense thanks so much :)